While trying to access SOHO's coronograph data we noticed that the latest LASCO C2 image is updated on August 18 at 13:36 and LASCO C3 at 13:30. Also, SOHO's EIT 304 latest image is from August 17 at 13:19, EIT 195 is from August 16 at 13:13 while EIT 171 and 284 show no data.
We checked up "keyhole" schedule but it is scheduled for September, not August.
Then, we start to look at SDO images but we came across similar problem. AIA and HMI images are not updated. Latest AIA image is from August 18 at 06:00 UTC as well as HMI. But this time we knew the "reason". It looks like air conditioning at Stanford University stopped at early hours on August 18. SDO twitted and announced on their website that HMI and AIA images will not be available until the problem is solved, possibly Monday. They stated that science data is unaffected but they don't have access to the computer drives that hold the images. Yeah, right...
No A.C. at Stanford, HMI & AIA Images Not Available
Sat, 18 Aug 2012
Late last night the air conditioning at Stanford University stopped. HMI & AIA images will not be available until the problem is solved, possibly Monday. The science data is unaffected, but we do not have access to the computer drives that hold the images.
STEREO Ahead and Behind EUVI, COR1 and COR 2 images were last updated on August 19 at 06:06 UTC. But then something got weird. Looking at STEREO Behind HI images we saw some kind of errors. Nothing like we saw before.
Then more errors. PROBA2 SWAP 174 data (at 13:45 UTC on August 19) and MLSO Mk4.
We looked at EVE imagery and found some black object left from the Sun. Kind of puzzled?!
Then all from the sudden we noticed strange light trails ahead of Sun - EVE and GOES 15
Well, something is going on with our sun observing instruments and we would like to get some answers.
If you are interested as well, stay tuned with us and come back for more updates!
THIS STORY WILL BE UPDATED so come back if you want to know more
It looks like SOHO LASCO coronographs are back again.
New message from SDO: Chilled water is still off at Stanford and HMI and AIA images remain unavailable. Best estimate is they will be available Monday afternoon.
Latest STEREO Behind COR 2 image from 18:55 UTC is showing some error.
SOHO's coronographs are back again. However, most of the images from the time when coronographs were down are with errors.
Enlil data shows that STEREO Ahead was hit by CME cloud.
Intensive solar flare took place off the farside of the Sun producing a bright CME (seen on STEREO Ahead COR2 images).
CHECK IT FOR YOURSELF: STEREO, SDO, SOHO, MLSO (HAO/NCAR), SIDC